• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name dhamilton

    Log entry time 07:59:01 on April10,2008

    Entry number 231141

    keyword=Owl Shift Summary

    Owl Shift Summary
    David Hamilton
    Brian Hahn
    Khamit Ardashev

    00:00 - shift started with 4 microamps on target at production kinematics C during run 2675.

    00:15 - reduced ps3 by a factor of 10 in order to give a T3 rate of 10Hz. Increaed beam current to 5 microamps, deadtime rose to 21 %. Beam current was redcued back to 4 microamps.

    00:35 - ROC5 crashed during run 2677. It was rebooted and we started run 2678.

    05:10 - We noticed deadtime climbing to > 50 % during runs 2699 and 2701.

    It seems to me that this problem could be netowrk related. For the second night in a row, we noticed that the increase in deadtime and ROC5 hanging up coincides with a temporary loss of connection to the BB HV portserver.

    05:45 - data-taking restored after restarting the BB HV gui several times, ROC5 reboot and CODA restart.

    06:15 - beam trip caused y position to come back way off. Tripped BB HV.

    08:00 Shift ended with 4 microamps on target.

    In excess of 34 million good events taken during the night.